A wheel has a constant angular acceleration of 5rad/sec^2 starting from rest it turns through 300 rad.
A, what is the final angular velocity?
B, how much time elapses while it turns through the 300 radians?
1
Expert's answer
2020-05-07T08:40:41-0400

(A) the final angular velocity

"\omega_f=\sqrt{\omega_i^2+2\alpha \theta}\\
=\sqrt{0^2+2\times 5 \times 300}=55\:\rm rad/s."

(B) the time that elapses while a wheel turns through the 300 radians

"t=\sqrt{\frac{2\theta}{\alpha}}
=\sqrt{\frac{2\times 300}{5}}=11\:\rm s."